OpenStudy (anonymous):

Find the radius of a circle whose circumference is 216 units, use 3.14 for π. r = 34.395 r = 68.790

OpenStudy (ipwnbunnies):

\[Circumference = 2\pi r\] Therefore, \[r = \frac{circumference}{2\pi}\]

OpenStudy (ipwnbunnies):

Where 'r' is the radius of the circle.

OpenStudy (solomonzelman):

you seem not to reply, is there anything that troubling you with this?